# How Does Perfect Fitting Affect Representation Learning? On the Training Dynamics of Representations in Deep Neural Networks

### v1

@@ Abstract
In this paper, we elucidate how representations in deep neural networks (DNNs) evolve during training. We focus on overparameterized learning settings where the training continues much after the trained DNN starts to perfectly fit its training data. We examine the evolution of learned representations along the entire training process, including its perfect fitting regime, and with respect to the epoch-wise double descent phenomenon. We explore the representational similarity of DNN layers, each layer with respect to its own representations throughout the training process. For this, we use two similarity metrics: (1) The centered kernel alignment (CKA) similarity; (2) Similarity of decision regions of linear classifier probes that we train for the DNN layers. Our extensive experiments discover training dynamics patterns that can emerge in layers depending on the relative layer-depth, DNN width, and architecture. We show that representations at the deeper layers evolve much more in the training when an epoch-wise double descent occurs. For Vision Transformer, we show that the perfect fitting threshold creates a transition in the evolution of representations across all the encoder blocks.
@@ 1 Introduction
Usually, DNNs are overparameterized models, namely, they have many more learnable parameters than training examples. Due to such overparameterization, DNNs are often trained to perfectly fit their training data (e.g., correctly labeling all the training examples in a classification task). This perfect fitting property stands in sharp contrast to conventional machine learning guidelines that associate overfitting (and, hence, perfect fitting) with poor generalization performance. Yet, DNNs that perfectly fit their training data often show good generalization performance, raising foundational questions on their inner workings.
The training process of overparameterized models can be divided into the following two parts. The first part is from the beginning of training (initialization) and until the time point (training epoch) when the model perfectly fits its training data for the first time. Then, the second part is the perfect fitting regime of training where the trained model always perfectly fits its training data (Fig. 1g). Evaluating the test error of the model as a function of the training epoch can show (especially when label noise exists) an epoch-wise double descent phenomenon [1] that includes two parts (Figs. 2g,7g): (i) A classical “U-shape” that reflects the classical approach of an ideal early stopping point: the test error first decreases, then increases back and peaks towards the start of the perfect fitting regime. (ii) A decreasing test error after its peak and into the entire perfect fitting regime that can eventually achieve a low test error, suggesting the approach of long training as an alternative to early stopping.
@@
In this paper, we ask how do overparameterization and perfect fitting affect the evolution of representations in each of the DNN layers during training? To address this question, we evaluate how similar are the representations that evolve at each layer along the entire training process. Namely, for a layer of interest, we provide a detailed map that shows how similar are the representations of that layer at every two different time points (epochs) in the training. We believe that our evaluations provide the most extensive and detailed elucidation of representational similarity in the training process (as we discuss in Section 2, some previous works had a very limited evaluation of representational similarity in training with respect to only a particular training time-point).
@@
In our evaluations, we use two different approaches to quantify representational similarity. First, the centered kernel alignment (CKA) [2], which is a popular measure of representational similarity between layers of different DNNs or within the same DNN. Second, we propose a new approach to measure representational similarity by using linear classifier probes [3] for the DNN layers during their training process. We use the linear classifier probes of the same layer at different training epochs and evaluate the similarity of their decision regions as a measure of similarity between the representations at these training epochs.
@@
Using our detailed representational similarity evaluation, we examine overparameterized learning settings where the training continues much after the DNN starts to perfectly fit its training data. By this, we elucidate the evolution of representations at the various parts of training, including the perfect fitting regime, and in settings with/without the epoch-wise double descent phenomenon. This sheds new light on representation learning in overparameterized models, including the following insights:
@@
    Representations at the first layer, in the perfect fitting regime of training a sufficiently wide DNN, are more similar to random representations (e.g., as at the initialized, untrained DNN) than to representations at the best early stopping epochs before perfect fitting starts.
@@
    Representations at the deeper layers evolve much more in the training when an epoch-wise double descent occurs. This is especially apparent in ResNet training.
@@
    Only the linear classifier probes of the deeper layers have an epoch-wise double descent (in their own, probing test error) when the entire DNN also has the epoch-wise double descent. This suggests that the representations and features at the deeper layers have a crucial role in creating the epoch-wise double descent phenomenon of the entire DNN.
@@
    Representations often have one or two transitions in their training evolution. These create training segments in which representations are relatively similar, visually appearing as lighter color “similarity blocks” along the main diagonal of the similarity diagrams. For ResNet, this is especially visible when label noise and epoch-wise double descent exist (see, e.g., Figs. 2c-2f). For Vision Transformer (ViT), this is well observed for both the noiseless and label noise settings (Figs. 3, 17); moreover, we show that the perfect fitting threshold creates a transition in the evolution of representations across all the ViT encoder blocks.
@@ 2 Related Work
The effect of DNN width on similarity of representations was examined in [4] using the CCA metric, and in [2, 5] using the CKA metric. In contrast to these works, our focus is on elucidating the representional similarity during the training process. In the following, we discuss the existing literature which is more related to our work.
@@ 2.1 Representational Similarity Throughout the Training Process
In [6], the SVCCA metric for representational similarity is proposed and utilized for several analyses, including showing how the representational similarity between the model during training and the model at the end of training evolves during training of a convolutional model and a ResNet model. In addition, [7] shows the evolution of the SVCCA metric during 4 epochs of a BERT model training. In this paper, we examine the evolution of representational similarity much more extensively than previous works, we consider various new settings (including for a ViT architecture, noiseless and noisy datasets) and, most importantly, we use it for the first time as a tool for analyzing the perfect fitting effect and the epoch-wise double descent phenomenon.
@@
In [8], the loss landscape (specifically, connectivity of local minima) and representational similarity (using the CCA metric [6]) are examined for various training heuristics from the common practice, such as learning rate restarts, warmup, and distillation. Their results include evaluation of CCA representational similarity between a VGG model at training iteration 0 and the specific training iteration at the end of their warmup phase (if applied); their evaluation reflects high similarity for the first convolutional layer of the model. They also evaluate the CCA similarity between the model at the end of training and at the intermediate training iteration after the warmup. Importantly, the work in [8] does not show or analyze the detailed evolution of the representational similarity during the entire training process. In contrast, in our paper here we examine the representational similarity in much more detail, enabling to understand the representational similarity of various pairs of epochs, throughout a long training process with a small constant learning rate that does not affect the natural training dynamics. Moreover, we focus here on understanding how the representational similarity evolves in overparameterized learning where perfect fitting does not prevent good generalization.
2.2 Representational Similarity between Trained and Untrained Models
In [2], the CKA similarity between trained and untrained convolutional DNNs was examined. They show that the representations in the first layers of the DNN are similar to those of the corresponding layers in the untrained DNN. Their CKA values for the first layer of a 10-layer convolutional model on CIFAR-10 was slightly above 0.8. In this paper (e.g., Figs. 1a, 2a, 7a), we show that the first layer can have CKA similarity (nearly) 1 in the perfect fitting regime of training.
In Appendix E we discuss the related work [9] and explain how we disprove the generality of their claim regarding the learning process of the first layer in the DNN. We also discuss in Appendix E an additional related work [10].
@@ 3.1 The CKA Measure: Definition and Formulation
Quantifying similarity of representations in intermediate DNN layers is known as a challenging task due to the inconsistent and intricate role the various neurons have in making the representations. The CKA was proposed in [2] as a robust method for measuring similarity of representations; indeed, the CKA (based on a linear kernel) is a popular, leading choice for analyzing representations in DNNs (e.g., [5, 11]).
@@
The CKA is computed for two layers of possibly different DNNs, namely, $f_{L}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$ and $g_{L^{\prime}}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L^{\prime}}}$ based on a set of $m$ inputs examples $\mathcal{S}_{m}=\{\mathbf{x}_{1},...,\mathbf{x}_{m}\}\subset\mathcal{X}$. For these examples, the representations are computed and organized in the following matrices: $\mathbf{F}\triangleq\left[f_{L}\left(\mathbf{x}_{1}\right),...,f_{L}\left(\mathbf{x}_{m}\right)\right]^{T}$ and $\mathbf{G}\triangleq\left[g_{L^{\prime}}\left(\mathbf{x}_{1}\right),...,g_{L^{\prime}}\left(\mathbf{x}_{m}\right)\right]^{T}$. Then, $\mathbf{K}\triangleq\mathbf{F}\mathbf{F}^{T}$ and $\mathbf{L}\triangleq\mathbf{G}\mathbf{G}^{T}$ are the Gram matrices of the representations (separately for $f_{L}$, $g_{L^{\prime}}$) whose components quantify the representation similarity between pairs of examples. These Gram matrices further go through centering of their rows and columns, namely, computing $\mathbf{K}^{\prime}=\mathbf{H}\mathbf{K}\mathbf{H}$ and $\mathbf{L}^{\prime}=\mathbf{H}\mathbf{L}\mathbf{H}$ where $\mathbf{H}\triangleq\mathbf{I}_{m\times m}-\frac{1}{m}\mathbf{1}_{m\times m}$ is the centering matrix, $\mathbf{I}_{m\times m}$ is the $m\times m$ identity matrix, and $\mathbf{1}_{m\times m}$ is an $m\times m$ matrix of ones. Then, the HSIC is computed as the normalized dot product of the vectorized forms of $\mathbf{K}^{\prime}$, $\mathbf{L}^{\prime}$:
${\sf HSIC}_{0}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})=\frac{1}{m-1}{\sf vec}\left(\mathbf{K}^{\prime}\right)\cdot{\sf vec}\left(\mathbf{L}^{\prime}\right)$ (1)
where $\mathbf{K}^{\prime}$ and $\mathbf{L}^{\prime}$ were defined above based on $f_{L}$ and $g_{L^{\prime}}$, respectively, and the set of examples $\mathcal{S}_{m}$. Importantly, the HSIC measure is not affected by permutation of neurons and orthogonal transformations of the representations. The important property of invariance to uniform scaling of the representations is added by the following normalization that defines the CKA measure:
${\sf CKA}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})=\frac{{\sf HSIC}_{0}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})}{\sqrt{{\sf HSIC}_{0}(f_{L},f_{L};\mathcal{S}_{m})}\sqrt{{\sf HSIC}_{0}(g_{L^{\prime}},g_{L^{\prime}};\mathcal{S}_{m})}}.$ (2)
3.2 The Proposed CKA Evaluation During the Training Process
@@
Figure 1: CKA evaluations for ResNet-18 (of standard width), trained on CIFAR-10 without label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows the test and train errors during training.
Figure 2: CKA evaluations for ResNet-18 (standard width), trained on CIFAR-10 with 20% label noise.
@@
In this section we examine the evolving representations of the DNN layers throughout the training. The similarity of representations in each layer are computed with respect to the representations of the same layer, at different time points (epochs) in the training process. This can be formulated as follows. The mapping from the input to the representation of the $L^{\sf th}$ layer at training epoch $t\in\{0,\dots,3999\}$ is denoted as $f_{L}^{(t)}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$, where $p_{L}$ is the dimension of the representation of layer $L$. Then, we compute the CKA for $f_{L}^{(t)}$ and $f_{L}^{(t^{\prime})}$ for every two training epochs $t,t^{\prime}\in\{0,\dots,3999\}$ [^1] In experiments on ResNet and CIFAR-10, we always evaluated the entire detailed epoch grid from epoch 0 to 3999. In the other experiments, we downsampled the epoch grid for which the CKA is computed, due to the high computational and storage demands; even then, our epoch grid is relatively dense w.r.t. the presentation of the results using a logarithmic epoch axis. The details on the epoch grids are provided in Appendix F.. Namely, for a layer $L$, we create a 2D diagram (heatmap) of CKA values, ${\sf CKA}(f_{L}^{(t)},f_{L}^{(t^{\prime})};\mathcal{S}_{m})$ for $t,t^{\prime}\in\{0,\dots,3999\}$.
@@
We trained ResNet-18 and ViT-B/16 DNNs on the CIFAR-10 dataset (without and with 20% label noise). The ResNet-18 was trained also on the Tiny ImageNet dataset. Unless otherwise specified, models were trained for 4000 epochs with the Adam optimizer at a constant learning rate of 0.0001 (without weight decay) that ensures smooth and natural evolution of the training process (this was also the leading setting in the deep double descent study in [1]). In addition, we provide exemplary results for the SGD optimizer with/without momentum (Appendix Figs. 13, 14, 15), and for Adam with weight decay (see Appendix Fig. 16). Standard data augmentation was applied. Random initialization was applied, epoch 0 corresponds to the initialized DNN before its training starts. See Appendix G for the summary of experiment settings.
@@
The resulting CKA diagrams are provided in Figs. 1, 2, 3 and in Appendix Figs. 7-17. Note that the CKA of two training epochs is symmetric, ${\sf CKA}(f_{L}^{(t)},f_{L}^{(t^{\prime})};\mathcal{S}_{m})={\sf CKA}(f_{L}^{(t^{\prime})},f_{L}^{(t)};\mathcal{S}_{m})$. Naturally, the similarity is maximal along the main diagonal that connects the lower-left and upper-right corners, as values along this diagonal correspond to similarity of a layer at a specific epoch with itself. Usually, the similarity is relatively high around the main diagonal, as representations usually evolve relatively smooth throughout the training process.
@@
Each of the 2D diagrams considers a specific layer in the DNN and can be read according to its rows. Each row is associated with a different training epoch to consider its representation as a reference to evaluate the similarity for the entire training process, which corresponds to the horizontal axis of the diagram. For example, the first (bottom) row of the diagram shows how the representations throughout the entire training process are similar to the random representations at the initialization. In the following, we discuss the various findings that stem from these CKA diagrams.

### v2

# How Do the Architecture and Optimizer Affect Representation Learning? On the Training Dynamics of Representations in Deep Neural Networks

@@ Abstract
In this paper, we elucidate how representations in deep neural networks (DNNs) evolve during training. Our focus is on overparameterized learning settings where the training continues much after the trained DNN starts to perfectly fit its training data. We examine the evolution of learned representations along the entire training process. We explore the representational similarity of DNN layers, each layer with respect to its own representations throughout the training process. For this, we use two similarity metrics: (1) The centered kernel alignment (CKA) similarity; (2) Similarity of decision regions of linear classifier probes that we train for the DNN layers. We visualize and analyze the decision regions of the DNN output and the layer probes during the DNN training to show how they geometrically evolve. Our extensive experiments discover training dynamics patterns that can emerge in layers depending on the relative layer-depth, architecture and optimizer. Among our findings: (i) The training phases, including those related to memorization, are more distinguishable in SGD training than in Adam training, and for Vision Transformer (ViT) than for ResNet; (ii) Unlike ResNet, the ViT layers have synchronized dynamics of representation learning.
Machine Learning, ICML
@@ 1 Introduction
Usually, DNNs are overparameterized models, namely, they have many more learnable parameters than training examples. Due to such overparameterization, classification DNNs are often trained to perfectly fit their training data, i.e., to correctly label all their training examples (according to the pairings in the given dataset). This perfect fitting property stands in sharp contrast to conventional machine learning guidelines that associate overfitting (and, hence, perfect fitting) with poor generalization performance. Yet, DNNs that perfectly fit their training data often show good generalization performance, raising foundational questions on their inner workings.
@@
In this paper, we ask how do the DNN architecture and optimization method affect the evolution of representations in various DNN layers during training? To address this question, we evaluate how similar are the representations that evolve at each layer along the entire training process. Namely, for a layer of interest, we provide a detailed map that shows how similar are the representations of that layer at every two different time points (epochs) in the training. We believe that our evaluations provide the most extensive and detailed elucidation of representational similarity in the training process (as we discuss in Section 2, some previous works had a very limited evaluation of representational similarity in training with respect to only a particular training time-point).
In our evaluations, we use two different approaches to quantify representational similarity:
  • The centered kernel alignment (CKA) (Kornblith et al., 2019), which is a popular measure of representational similarity between layers of different DNNs or within the same DNN.
@@
  • We propose a new approach to measure representational similarity by using linear classifier probes (Alain & Bengio, 2017) for the DNN layers during their training process. We use the linear classifier probes of the same layer at different training epochs and evaluate the similarity of their decision regions as a measure of similarity between the representations at these training epochs. Our layerwise decision regions similarity (DRS) approach allows interpretability by visualizing decision regions that participated in the similarity evaluation at different epochs (e.g., Figs. 4, 5).
@@
Using our detailed representational similarity evaluation, we examine overparameterized learning settings where the training continues much after the DNN starts to perfectly fit its training data. By this, we elucidate the evolution of representations at the principal phases of training:
  • Phase I: Learning of general representations that do not overfit.
  • Phase II: Refining representations in order to memorize mislabeled and/or atypical examples and achieving perfect fitting of the entire training dataset.
  • Phase III: Additional evolution of representations while maintaining the perfect fitting. This evolution might improve the representations but not necessarily.
Our representational similarity diagrams emphasize these three training phases as similarity blocks where, in each, the representations are relatively similar but differ from the other training phases (thus the similarity blocks visually appear as lighter color blocks along the main diagonal of the similarity diagrams, e.g., see Fig. 2). We further explore the training dynamics by the DRS approach and its visualization capability (e.g., Figs. 4, 5). All of these shed new light on representation learning in overparameterized models, including the following insights:
@@
    Training phases appear as similarity blocks in more of the DNN layers, and are more clearly identified (implying relatively-faster transition between phases and/or greater dissimilarity between representations of different training phases), in
    1. (a)
      ViT than in ResNet
    2. (b)
      SGD training than in Adam training
@@
    Training dynamics of representations are synchronized across the layers in ViT, but not in ResNet.
@@
    Representations at the first layer, in the perfect fitting regime of Adam training of a sufficiently wide DNN, are more similar to random representations (e.g., as at the initialized, untrained DNN) than to representations at the best early stopping epochs before perfect fitting starts. This phenomenon is consistently observed in Adam training of convolutional networks (such as ResNet or a two-layer model) but not for SGD and ViT. We show that the adaptive normalization of gradients by Adam is necessary for this phenomenon to occur.
@@
    The decision regions of the DNN output and the deeper-layer probes are likely to become more fragmented during the training phase of memorizing atypical examples (in noiseless experiments), and then the fragmentation level stabilizes in the perfect fitting training phase (possibly with some reduction in fragmentation before the stabilization). This behavior is observed more clearly for ViT, also due to its synchronized training dynamics across its layers, showing that the implicit regularization of the optimizer during training can significantly depend on the DNN architecture.
@@ 2.1 Representational Similarity Throughout the Training Process
Raghu et al. (2017) proposed the SVCCA metric for representational similarity and used it for several analyses, including showing how the representational similarity between the model during training and the model at the end of training evolves during training of a convolutional model and a ResNet model. Liu et al. (2021) showed the evolution of the SVCCA metric during 4 epochs of a BERT model training. Gotmare et al. (2019) analyze CCA representational similarity among few training time-points in a VGG model training (see more details in Appendix A.1), but they do not show or analyze the detailed evolution of the representational similarity during the entire training process. In this paper, we examine the evolution of representational similarity much more extensively than previous works, we consider various new settings (such as for a ViT architecture). In contrast to previous works, in our paper here we examine the representational similarity in much more detail, enabling to understand the representational similarity of various pairs of epochs, throughout a long training process with a small constant learning rate that does not affect the natural training dynamics. Moreover, we focus here on understanding how the representational similarity evolves in overparameterized learning where perfect fitting does not prevent good generalization.
In Appendix A.2, we discuss the related works by Kornblith et al. (2019); Chowers & Weiss (2023); Hermann & Lampinen (2020) who studied the similarity between trained and untrained models.
2.2 Training Phases and Memorization
Zhang et al. (2017) showed that DNNs have the ability to perfectly fit (i.e., memorize) pure noise. Arpit et al. (2017) showed that the training process learns real data patterns before memorizing the noise; and that even without artificially adding noise to real data, the training examples can vary from easy to hard in their learning difficulty – unlike in pure noise where all examples are similarly hard to learn. Liu et al. (2020) showed that learning of classifiers from training data with label noise has two phases. First, an early phase where the training examples with correct labels dominate the learning (due to dominating the overall loss gradient), which yields a relatively high prediction accuracy also for the training examples with the wrong labels (this prediction accuracy is according to the true labels that are unknown in the training for these examples). Then, the next phase memorizes the wrong labels, enabling the model to perfectly fit the training data at the expense of the high prediction accuracy (w.r.t. the true labels) it had earlier in training.
Maini et al. (2023) showed that memorization of mislabeled or atypical training examples (identified by a method by Jiang et al. (2021)) is usually not limited to the last nor other specific layers of the DNN. They showed that such memorization can be induced by a small set of neurons dispersed across the DNN layers, and that this neuron set can vary due to learning problem, training procedure, DNN architecture, etc. They based their claims using layer-wise gradient evaluation, rewinding/retraining individual layers, and optimization-based search for neurons critical for memorization.
Differently from these previous works, we elucidate the training phases and memorization from the perspective of representations at intermediate layers of the DNNs.
2.3 Inductive Bias of DNN Architectures
Raghu et al. (2021) analyzed the representational similarity and differences between ResNet and ViT fully trained models, but not during their training like we do here. Somepalli et al. (2022) studied the inductive bias of various DNNs by evaluating similarity between decision regions of the prediction output of fully-trained DNNs, measuring their fragmentation level, and visualizing them. Specifically, they showed that convolutional architectures have less fragmented decision regions compared to architectures such as ViT. We significantly extend their analysis technique and insights, as we here study decision regions’ evolution during the training process and for intermediate layers using probes.
2.4 The Implicit Regularization of the Adam Optimizer
Kunstner et al. (2023); Cattaneo et al. (2024) analyzed the implicit regularization of the Adam optimizer, which significantly differs from SGD due to the adaptive gradient normalization in Adam. We discuss the work by Cattaneo et al. (2024) in more detail in Section 3.3.3 together with our results that contribute to better understanding of the implicit regularization of Adam on the representation learning dynamics.
@@ 3.1 The CKA Measure: Definition and Formulation
Quantifying similarity of representations in intermediate DNN layers is known as a challenging task due to the inconsistent and intricate role the various neurons have in making the representations. The CKA was proposed by Kornblith et al. (2019) as a robust method for measuring similarity of representations; indeed, the CKA (based on a linear kernel) is a popular, leading choice for analyzing representations in DNNs (e.g., Nguyen et al., 2021; Raghu et al., 2021).
@@
The CKA is computed for two layers of possibly different DNNs, namely, $f_{L}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$ and $g_{L^{\prime}}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L^{\prime}}}$ based on a set of $m$ inputs examples $\mathcal{S}_{m}=\{\mathbf{x}_{1},...,\mathbf{x}_{m}\}\subset\mathcal{X}$. For these examples, the representations are computed and organized in the following matrices: $\mathbf{F}\triangleq\left[f_{L}\left(\mathbf{x}_{1}\right),...,f_{L}\left(\mathbf{x}_{m}\right)\right]^{T}$ and $\mathbf{G}\triangleq\left[g_{L^{\prime}}\left(\mathbf{x}_{1}\right),...,g_{L^{\prime}}\left(\mathbf{x}_{m}\right)\right]^{T}$. Then, $\mathbf{K}\triangleq\mathbf{F}\mathbf{F}^{T}$ and $\mathbf{L}\triangleq\mathbf{G}\mathbf{G}^{T}$ are the Gram matrices of the representations (separately for $f_{L}$, $g_{L^{\prime}}$) whose components quantify the representation similarity between pairs of examples. These Gram matrices further go through centering of their rows and columns, namely, computing $\mathbf{K}^{\prime}=\mathbf{H}\mathbf{K}\mathbf{H}$ and $\mathbf{L}^{\prime}=\mathbf{H}\mathbf{L}\mathbf{H}$ where $\mathbf{H}\triangleq\mathbf{I}_{m\times m}-\frac{1}{m}\mathbf{1}_{m\times m}$ is the centering matrix, $\mathbf{I}_{m\times m}$ is the $m\times m$ identity matrix, and $\mathbf{1}_{m\times m}$ is an $m\times m$ matrix of ones. Then, the HSIC is computed as the normalized dot product of the vectorized forms of $\mathbf{K}^{\prime}$, $\mathbf{L}^{\prime}$: ${\sf HSIC}_{0}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})={\sf vec}\left(\mathbf{K}^{\prime}\right)\cdot{\sf vec}\left(\mathbf{L}^{\prime}\right)/{(m-1)}$ where $\mathbf{K}^{\prime}$ and $\mathbf{L}^{\prime}$ were defined above based on $f_{L}$ and $g_{L^{\prime}}$, respectively, and the set of examples $\mathcal{S}_{m}$. Importantly, the HSIC measure is not affected by permutation of neurons and orthogonal transformations of the representations. The important property of invariance to uniform scaling of the representations is added by the following normalization that defines the CKA measure:
${\sf CKA}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})=$ (1)
$~{}~{}\frac{{\sf HSIC}_{0}(f_{L},g_{L^{\prime}};\mathcal{S}_{m})}{\sqrt{{\sf HSIC}_{0}(f_{L},f_{L};\mathcal{S}_{m})}\sqrt{{\sf HSIC}_{0}(g_{L^{\prime}},g_{L^{\prime}};\mathcal{S}_{m})}}.$
3.2 The Proposed CKA Evaluation During Training
@@
Figure 1: CKA evaluations for ResNet-18, trained using Adam on CIFAR-10 without label noise. Each of the (a)-(f) subfigures shows the CKA representational similarity of a specific layer in the ResNet-18 during its training. ResNet Block $j$ refers to the representation at the end of the $j^{\sf th}$ block of layers in the ResNet-18 architecture. (g) shows train (of all the train dataset), atypical train examples, and test errors during training. Dashed lines denote transition from training phase I to II (gray) and training phase II to III (cyan).
@@
In this section we examine the evolving representations of the DNN layers throughout the training. The similarity of representations in each layer are computed with respect to the representations of the same layer, at different time points (epochs) in the training process. This can be formulated as follows. The mapping from the input to the representation of the $L^{\sf th}$ layer at training epoch $t\in\{0,\dots,3999\}$ is denoted as $f_{L}^{(t)}:\mathcal{X}\rightarrow\mathbb{R}^{p_{L}}$, where $p_{L}$ is the dimension of the representation of layer $L$. Then, we compute the CKA for $f_{L}^{(t)}$ and $f_{L}^{(t^{\prime})}$ for every two training epochs $t,t^{\prime}\in\mathcal{T}\subset\{0,\dots,3999\}$, where $\mathcal{T}$ is a relatively dense subset of training epochs from the entire training process (we did this grid downsampling due to the high computational and storage demands, and even then, our epoch grid is relatively dense w.r.t. the presentation of the results using a logarithmic epoch axis; the details on the epoch grids are provided in Appendix B). Namely, for a layer $L$, we create a 2D diagram (heatmap) of CKA values, ${\sf CKA}(f_{L}^{(t)},f_{L}^{(t^{\prime})};\mathcal{S}_{m})$ for $t,t^{\prime}\in\{0,\dots,3999\}$.
@@
Our experiments mainly consider ResNet-18 and ViT-B/16 DNNs, training using Adam and SGD, and CIFAR-10 (without and with 20% label noise) and SVHN datasets. The ResNet-18 was trained also on the Tiny ImageNet dataset. Unless otherwise specified, models were trained for 4000 epochs with the Adam optimizer at a constant learning rate of 0.0001 (without weight decay) that ensures smooth and natural evolution of the training process (this was also the leading setting in the deep double descent study by Nakkiran et al. (2021)). We show results for SGD with/without momentum (e.g., Figs. D.8, D.9, D.10), and for Adam with weight decay (Fig. D.11). Standard data augmentation was applied. Random initialization was applied, epoch 0 corresponds to the initialized DNN before its training starts. See Appendix C and Table C.1 for the summary of all experiment settings.
@@
The resulting CKA diagrams are provided in Figs. 1, 2 and in Appendix Figs. D.1-D.24. Note that the CKA of two training epochs, from the same training process, is symmetric, ${\sf CKA}(f_{L}^{(t)},f_{L}^{(t^{\prime})};\mathcal{S}_{m})={\sf CKA}(f_{L}^{(t^{\prime})},f_{L}^{(t)};\mathcal{S}_{m})$. Naturally, the similarity is maximal along the main diagonal that connects the lower-left and upper-right corners, as values along this diagonal correspond to similarity of a layer at a specific epoch with itself. Usually, the similarity is relatively high around the main diagonal, as representations usually evolve relatively smooth throughout the training process.
@@
Each of the 2D diagrams is for a specific layer in the DNN and can be read according to its rows. Each row is associated with a different training epoch to consider its representation as a reference to evaluate the similarity for the entire training process, which corresponds to the horizontal axis of the diagram. E.g., the first (bottom) row of the diagram shows how the representations throughout the entire training process are similar to the random representations at the initialization. With the CKA diagrams, we provide error graphs of test data (in yellow line), train data (all the train data, in black line), and the atypical train examples (in green line, available for selected experiments without artificially added label noise). We explain in Appendix B.4 how the atypical train examples were found.
The dashed lines denote the (estimated) transition points between the training phases. The gray dashed line denotes the transition from Phase I of learning general features to Phase II of memorizing mislabeled/atypical examples; this transition point occurs towards the arrival of the mislabeled/atypical error graph to its steepest decrease slope (due to the intricacy of these real training dynamics, the exact transition point between Phase I and Phase II does not follow a clear mathematical condition for all the experiments and therefore we manually annotate it according to the CKA diagrams). The cyan dashed line denotes when perfect fitting of the entire training dataset starts, namely, the transition to Phase III of training; we estimate this transition point as the epoch at which the train error (of the entire dataset) starts to be below 0.001 (because accurate perfect fitting of zero train error is rare for real datasets and iterative training using mini-batches; note that the beginning of the similarity block may not necessarily exactly align with the arrival to zero train error, this behavior agrees with the deep double descent analysis by Nakkiran et al. (2021) where the test error peak is not necessarily exactly aligned with the arrival to zero train error).
In the following, we discuss the various findings that stem from these CKA diagrams.